### v1

# Generalizable 3D Scene Reconstruction via Divide and Conquer from a Single View

Abstract
Single-view 3D reconstruction is currently approached from two dominant perspectives: reconstruction of scenes with limited diversity using 3D data supervision or reconstruction of diverse singular objects using large image priors. However, real-world scenarios are far more complex and exceed the capabilities of these methods. We therefore propose a hybrid method following a divide-and-conquer strategy. We first process the scene holistically, extracting depth and semantic information, and then leverage a single-shot object-level method for the detailed reconstruction of individual components. By following a compositional processing approach, the overall framework achieves full reconstruction of complex 3D scenes from a single image. We purposely design our pipeline to be highly modular by carefully integrating specific procedures for each processing step, without requiring an end-to-end training of the whole system. This enables the pipeline to naturally improve as future methods can replace the individual modules. We demonstrate the reconstruction performance of our approach on both synthetic and real-world scenes, comparing favorable against prior works. Project page: https://andreeadogaru.github.io/Gen3DSR.
Keywords:
3D scene reconstruction single-view compositional
1 Introduction
Refer to caption Refer to caption Refer to caption Refer to caption
Input image 3D Scene reconstruction 3D Scene editing
Scene components
Figure 1: Our method can reconstruct a scene in 3D from a single input view. The scene is semantically processed and partially occluded objects are completed. The individually reconstructed object instances are composed back into the scene using monocular depth guides. We represent each component with a triangle mesh, enabling downstream applications such as scene manipulation and editing.
Single-view 3D scene reconstruction refers to the problem of understanding and explaining all the visible components that assembled together create a 3D scene which closely reproduces the original 2D observation. The computer vision and graphics communities have long been interested in automating this task, yet its complexity still leaves room for many improvements [59, 30]. Successful single-view applications have been developed for specific purposes such as face reconstruction [14, 29], hair modeling [63], and many more. However, the 3D understanding from a single image task is far from solved in the case of larger scale problems such as indoor/outdoor scene reconstruction with multiple objects [54].
In general, even reconstructing one 3D object from a single image is a severely ill-posed problem, e.g. it is impossible to tell precisely how the back side of an object looks like if the input image only observes the front. Nonetheless, if the distribution of objects that are naturally present in our day-to-day lives is known, one can plausibly predict the shape and appearance of a 3D object from very limited information. Accordingly, various priors have been used in the context of particular object classes (such as simple shapes [1], or human faces [14]). However, modeling entire scenes is a significantly more challenging problem.
Given the complexity of real-world scenes, reversing the process of image capturing in an end-to-end fashion would require a huge amount of data covering the variability of realistic environments. Therefore, many works solve a simplified version of the task by focusing on single objects or indoor rooms with a limited number of objects in which they predict the scene layouts containing 3D instance bounding boxes and optionally object meshes. Under these assumptions, most of the existent solutions rely on 3D scene geometry supervision from synthetic datasets. This class of methods usually struggles when applied to real-world images due to the domain gap. In contrast, we propose to tackle the single-view 3D scene reconstruction problem in a divide-and-conquer approach while building on the research advancement of related, simpler, tasks. In figure 1 we show that, following this approach, our pipeline is able to reconstruct multi-object scenes with unprecedented quality.
In the past few years, the field of computer vision has seen tremendous progress in solving particular tasks such as depth estimation and single-image 3D object reconstruction. It is the right time for these components to be assembled to solve the challenging task of full 3D scene reconstruction. We have identified the following sub-problems that together comprehensively explain a 3D scene and enable its reconstruction from a single input image: estimating the camera calibration, predicting the (metric) depth map, segmenting entities, detecting foreground instances, reconstructing the background, recovering the occluded parts of the individual objects (amodal completion), and reconstructing them. Designed as a disentangled framework, our method is open for incremental improvements and future enhanced components can be easily plugged in to boost the reconstruction performance of the entire system.
The main contributions of this paper are outlined as follows:
  • We build a compositional framework that comprehensively reconstructs a scene without requiring end-to-end training.
  • We develop a model for amodal completion and show how it can be used towards achieving full scene reconstruction.
  • We design the connecting links which enable integrating individually reconstructed 3D objects into the scene layout by exploiting single-view depth estimation.
  • We demonstrate through extensive quantitative and qualitative evaluations that the proposed modular approach exhibits generalizable full 3D scene reconstruction abilities.
2 Related Work
Starting with the pioneering work of Lawrence Roberts [51], the task of recovering 3D scene properties such as geometry, texture, and layout from a single 2D image has been extensively studied. Learning-based advancements have substantially progressed for this challenging task. However, existing methods still have limited understanding and struggle to faithfully reconstruct complex realistic scenes. In this section, we review the major lines of work and highlight the advantages of our approach that enable state-of-the-art results for realistic 3D scene reconstruction from a single view.
Feed-forward scene reconstruction.
The direct regression conditioned on the input image used in many computer vision tasks can also be employed for 3D scene reconstruction. Provided the availability of large scene collections, such systems are trained end-to-end using 3D supervision. Most such works [10, 71, 9] rely on an encoder-decoder architecture that takes as input the image and predicts voxel grids containing scene properties such as geometry, semantics and instance labels. These methods have the advantage that by predicting the scene layout jointly with the containing objects, the object poses are intrinsically correct. However, these solutions suffer from resolution limitations, require large 3D data collections for training [15, 16] and usually do not generalize well to real-world scenes. In contrast, we use a modular framework that does not require end-to-end training or 3D supervision.
Factorized scene reconstruction.
Following the formulation of Tulsiani et al. [60], the scene can be considered as composed of different factors including layout, object shape and pose that together make up a 3D scene representation. As this approach reconstructs the scene components individually, it is beneficial for downstream applications and the reconstruction system can be designed with specialized components for each type of factor. Methods following this paradigm initially reconstructed the objects in the scene as bounding boxes [20, 56, 13] or voxels [60, 36, 33], while later works use meshes [17, 45, 21, 34] or neural fields [37, 70, 69, 57]. Apart from the elementary factors of [60], more recent methods also include camera attributes [45, 21, 37, 20], textures [65, 8, 67], lightning [65, 67], or even material properties [65, 67]. Depending on the focus of the method, some works rely on a retrieval algorithm for identifying the object candidates [65, 21, 23, 2, 34] or regress their geometry [17, 37, 70, 57]. The initially predicted scene components are sometimes further fine-tuned using derivative-free optimization [23, 21, 7] or differentiable rendering [17, 67, 65, 69]. Our method also uses a compositional scene reconstruction framework. Particularly in our pipeline, each component is trained individually on extensive datasets. This strategy allows our proposed solution to transcend predefined (limited) classes of objects and reconstruct diverse scenes effectively.
Single-view scene understanding.
The reconstruction of simple scenes composed of a single isolated object has seen numerous approaches along the years [18]. Nonetheless, enabled by the versatile diffusion-based models capable of generating realistic images [53, 47] and large collections of 3D objects [12, 11], there has recently been a surge of single-view 3D reconstruction methods [39, 38, 40, 58] capable of creating digital objects with unprecedented quality. Closely related to 3D scene reconstruction is monocular depth estimation [42], which predicts from a single image a 2.5D representation of the visible content. Recent methods focus on improving the generalizability of the estimator by following a training protocol employing large datasets [4, 66] or by fine-tuning a large image prior such as Stable Diffusion [53] for the depth estimation task [27]. A scene is also described by the camera used to capture the image [26, 35, 19] and by the separation between different instances [31, 49, 68]. Though the current state-of-the-art methods for each individual sub-task achieve robust results for real images, they only offer an incomplete explanation of the scene. Therefore, we compose them in an open framework capable of fully reconstructing complex real-world scenes from a single image.
3 Method
Figure 2: Method Overview: the input image is first analyzed collectively by an ensemble of state-of-the-art monocular models. Subsequently, the identified instances are individually processed, as elaborated in figure 3. The reconstructed objects, along with the modeled background, are composed into the final scene, rendered here and suitable to various applications.
Our method, as illustrated in figure 2, takes as input a single RGB image $I$ and predicts the full 3D scene reconstruction $R(I)$ represented as a collection of triangle meshes. The proposed solution does not require end-to-end training and instead relies on off-the-shelf models carefully integrated into a seamless framework. First, we parse the image of the scene by finding the composing entities, and estimating the depth and camera parameters. Then, we separate the identified entities in stuff (amorphus shapes) and things (characteristic shapes) [6]. To recover the full view of each object, we perform amodal completion on the masked crops of the instances. Each object is reconstructed individually in a normalized space and aligned to the view space using the scene layout guides from the depth map. Importantly, we address the differences in focal length, principal point, and camera-to-object distance between the two spaces through reprojection. Finally, we model the background as the surface that approximates the stuff entities collectively.
3.1 Scene analysis
Our framework decouples the object-agnostic from the object-specific processing, pursuing a balance between the representational power and the generalization capabilities of the composing models. The image characteristics that can be best explained by perceiving the scene as a whole are the perspective properties and the depth. Additionally, the global context provides valuable guides for the recognition of the objects.
We rely on the Perspective Fields [26] formulation for estimating the camera calibration, $K_{img}$. Following the per-pixel parametrization using up-vectors and latitude, the representation can accommodate arbitrary camera models. In our pipeline we employ conventional camera parameters, i.e. field-of-view and principal point, which can be directly estimated from the Perspective Fields.
Monocular depth estimation is essential to our approach as it provides geometrical clues of scene layout, enabling the reconstruction of complex scenes that go beyond a single object. The performance of a depth estimator network is mostly determined by its capacity and training data. As we aim for a generalizable system, the single-image depth estimation method should perform well for a wide variety of natural images.
To fully explain the image, we employ entity segmentation [49, 50], which segments all semantically-meaningful entities without being constrained to a predefined set of classes. The unrestricted segmentation model yields robust dense predictions suitable for real-world scenarios. This step partitions each image $I$ into instances $\{M_{i}\}$ that can be individually reconstructed to compose the whole scene. Furthermore, we consider the natural separation of the instances in thing and stuff using a universal image segmentation model. This enables our method to tailor the reconstruction process for each group, leveraging their unique properties, as outlined in 3.2 and 3.3 respectively. To enhance downstream applications and text-conditioned reconstruction models, we harness the entire scene context to recognize and label the previously identified entities. Each segmented entity $M_{i}$ serves as a prompt for the OVSAM model [68], which, empowered by the recognition capabilities of the CLIP foundation model [25], outputs a corresponding instance label $L_{i}$.
3.2 Instance processing
Figure 3: Detailed Overview over the processing steps of each instance. An image is processed through the scene analysis part of our framework as described in figure 2. We add an object recognition information for diffusion guided completion to restore partially occluded objects. Then, we perform reconstruction and align the result back to the input view space.
We consider an object $O_{i}$ to be a well-defined shape categorised as thing, identified by its entity mask $M_{i}$, with corresponding RGB-D region $I_{i}=I[M_{i}]$, $D_{i}=D[M_{i}]$, and a label $L_{i}$. The processing steps are illustrated in figure 3. We choose to reconstruct each instance individually in order to fully benefit from a view-conditioned 2D diffusion model $\mathcal{Z}$ [40, 11], trained using multi-view images of objects from large scale object collections [12, 11]. As the images used to train these models were rendered using a fixed predefined camera configuration $K_{crop}$, they do not generalize well to in-the-wild object crops. Therefore, we propose to reproject the object-associated pixels from the input image to the image space expected by the diffusion model.
To this end, we employ the previously estimated camera calibration $K_{img}$ and depth map $D_{i}$ to unproject the pixels into the view space as $P_{i}^{view}\in\mathbb{R}^{3}$. We then find the virtual camera that together with the desired intrinsics $K_{crop}$ closely match the observed image. Finally, we use the transformation to reproject the pixels to a crop $C_{i}$ that resembles the training domain of $\mathcal{Z}$.
For simple scenes in which there is no occlusion between instances, the crop $C_{i}$ obtained using the equation above represents the full view of the object $O_{i}$. However, this is not the case for most real-world scenes, and directly feeding $C_{i}$ to the object reconstruction method $\mathcal{R}$ would result in an incomplete object. Therefore, we propose to recover the missing parts of $C_{i}$ by leveraging the image prior embodied by pre-trained large-scale diffusion models like Stable Diffusion. We approach the task named amodal completion which deals with recovering the shape and appearance of partially visible instances as an image-to-image translation problem. Specifically, we train a model to predict the view $\hat{C}_{i}$ of the full object $O_{i}$ conditioned on the object parts depicted in $C_{i}$ and the label $L_{i}$. Given the difficulty of collecting well-segmented training images guaranteed to contain complete objects, we generate a new synthetic dataset specifically for this task. We render synthetic objects from a large collection and then obtain the conditioning images by masking out parts of the object with a randomly overlaying silhouette of another object. We use this dataset to fine-tune Stable Diffusion, following the methodology of InstructPix2Pix [5], and concatenate the encoded conditioning image to the noisy latent. The weights added to the base network are initialized with zero, while the rest are taken from the pre-trained model to benefit from the learned image prior.
The complete crop $\hat{C}_{i}$ can now be used as input for the single-image 3D object reconstruction method $\mathcal{R}$. As our pipeline is not restricted to a specific model, we describe their common structure. Using a view-conditioned diffusion prior such as Zero-1-to-3 [40, 11] enables these models to reconstruct a wide-range of objects from a single view without the need for 3D training supervision. The object is reconstructed using differentiable 3D representations, e.g. neural fields [43, 44], 3D Gaussians [28], that are either directly fitted to multi-view images generated by the diffusion prior, or by optimizing a Score Distillation Sampling-based loss [48] against the diffusion prior.
Then, a polygonal mesh $R_{i}^{obj}$ aligned to the input crop $\hat{C}_{i}$ is extracted from the neural field representation for use in downstream applications. The obtained mesh can optionally be further fine-tuned to refine the texture of the reconstructed object.
The reconstructed instance in the object space $R_{i}^{obj}$ is transformed to the view space with respect to the inverse transformation determined by the virtual camera used for projection. The obtained mesh $R_{i}^{view}$ is aligned to the object points $P_{i}$ up to an unknown scale factor $s_{i}$; that is because $\mathcal{R}$ reconstructs objects at an arbitrary scale. We estimate $s_{i}$ as the scale factor that minimizes the distance between the visible points in $R_{i}^{view}$ and $P_{i}^{view}$.
3.3 Background modeling
Compared to the objects contained in a scene, the entities categorised as stuff, e.g. walls or ceiling, usually have a simpler geometry, which can be partially approximated by the corresponding regions in the depth map. Nonetheless, large areas are occluded by the foreground objects. Though we could follow a similar amodal completion strategy as the one described in 3.2, we opt for a more straightforward approach that interpolates the known data points. We consider all the background instances as one described by their mask union $M_{bg}$ and the corresponding image $I_{bg}$ and depth data $D_{bg}$. Then, we fit one small Multi-layer Perceptron $f:\mathbb{R}^{3}\to\mathbb{R}$ to represent the signed distance function (SDF) to the unprojected background points $P_{bg}^{view}$, and another one $c:\mathbb{R}^{3}\to\mathbb{R}^{3}$ to model the color associated to them. Lastly, we sample a dense grid of points in the camera frustum and evaluate their SDF values. The background surface is defined as $\mathcal{S}=\{x\in\mathbb{R}^{3}|f(x)=0\}$ and can be extracted from the grid using marching cubes [41].
4 Experiments
4.1 Implementation details
As the proposed framework is not constrained to specific modules, we leverage the significant progress made by the computer vision community in the recent years towards solving the different sub-tasks which we describe in the method section.
In the scene analysis stage, we rely on CropFormer [53] for entity segmentation, OneFormer [24] for separating them in foreground instances and background entities, and Perspective Fields [26] for estimating the camera calibration. We mainly use Marigold [27] for depth estimation. However, the model predicts affine-invariant depth which differs by an unknown image-level offset and scale from the absolute physical units. During evaluations, we estimate these factors based on the ground truth depth available in the datasets to ensure that the reconstructions align with the target. For in-the-wild predictions, we empirically found that estimating the two unknowns of Marigold output based on a metric depth estimation, in our case, DepthAnything [66], achieves better results than using the latter by itself.
In the instance processing stage, we perform amodal completion on the reprojected crops using our model obtained by fine-tuning Stable Diffusion v1.5 [53]. We also sample several points in the instance’s mask and feed them to OVSAM[68] together with the input image to obtain the text prompt for guiding the diffusion. The completed object is then reconstructed using DreamGaussian[58]. As the model performs best with known camera elevation of the input view, we estimate it following the method introduced in [39]. Then, we find the 3D points of the reconstruction that correspond to the unprojected instance points, which serve as our layout guide, and compute the scale which aligns them.
4.2 Results
Method 3D-supervision-free Zero-shot Compositional
InstPIFu [37]
USL [17]
BUOL [9]
Uni-3D [71]
DreamGaussian [58]
DreamGaussian
w/ our reprojection
Ours
Table 1: As opposed to existing approaches, our method is at the same time compositional, able to generalize in a zero-shot manner, and does not require 3D-supervision for training.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input image BUOL [9] Uni-3D [71] InstPIFu [37] Ours
Figure 4: Qualitative results on the 3D-FRONT dataset [15]. The methods consider reconstruct full scenes including background regions and foreground instances.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Input image InstPIFu [37] DreamG[58] Ours GT geometry
Figure 5: Qualitative results on the 3D-FRONT [15] dataset illustrated under the input view and a second one chosen to highlight the reconstruction performance. For completeness, we also include the background in the original view, where it is available.
3D-FRONT [15] HOPE-Image [61]
Method Chamfer $\downarrow$ F-Score $\uparrow$ Chamfer $\downarrow$ F-Score $\uparrow$
InstPIFu [37] 0.124 74.14 - -
DreamGaussian [58] 0.207 41.09 3.038 29.57
DreamGaussian
w/ our reprojection
0.187 49.98 4.059 30.62
Ours 0.120 68.82 1.446 54.82
Table 2: Quantitative results on the reconstruction of foreground instances. On 3D-FRONT, we obtain comparable scores to InstPIFu which is trained supervised with 3D data on this dataset. No scores are reported for InstPIFu on the HOPE-Image dataset, as the method cannot recognize and reconstruct any of the objects included in the evaluation. On both datasets, we outperform the non-compositional approach, DreamGaussian.
We showcase the performance of our method using several datasets across diverse scenarios. For numerical evaluation we first consider 3D-FRONT [15, 16], a synthetic dataset of indoor rooms with available ground truth geometry. Due to the large scale of dataset, we manually sample 100 images from the test split of [37], avoiding the images with heavy scene occlusions (e.g., camera positioned behind a plant), unrealistic object intersections and scenes with very few objects. In addition, we use the 10 validation images of HOPE-Image [61] dataset containing household objects captured under two scenarios. Though the dataset is intended for object pose evaluation, we find the ground-truth object alignment to match the input images well-enough for our purpose.
We report the quantitative results using the widely-employed metrics for 3D reconstruction: Chamfer Distance [3] and F-Score [32]. Both are computed between densely sampled sets of points from the reconstructed meshes and the ground truth respectively. As we focus the evaluation on whole scenes, the points are uniformly sampled from the entire geometry of a scene. Therefore, larger objects impact the overall score more.
We compare our method against several solutions covering multiple approaches for 3D scene reconstruction. An overview of their capabilities is presented in table 1. BUOL[9] and Uni-3D [71] are both feed-forward scene reconstruction methods that have been trained with 3D supervision on the 3D-FRONT dataset. However, their training views are rendered in a simpler way compared to the renderings we use in our evaluation. Therefore, the poor performance of these methods evaluated numerically in table 3 and qualitatively in figure 4 shows their lack of generalization.
Method Chamfer $\downarrow$ F-Score $\uparrow$
InstPIFu [37] 0.119 70.63
BUOL [9] 0.294 37.04
Uni-3D [71] 0.448 32.97
Ours 0.099 75.33
Table 3: Quantitative evaluation on full scene reconstruction including background and foreground elements on the 3D-FRONT dataset [15]. Our method, which performs in a zero-shot fashion, outperforms all the baselines, which have been trained using 3D supervision on a version of this dataset.
InstPIFu [37] and USL [17] are both compositional approaches that rely on 2D and 3D object detectors for identifying the objects to be reconstructed and aligning them in the scene. This limits their reconstruction capabilities to a small number of classes. Being trained on similar scenes, InstPIFu has a significant advantage in our evaluation on the 3D-FRONT dataset. Nevertheless, our method surpasses its quantitative performance on full scene reconstruction (table 3) and closely matches the one on foreground reconstruction (table 2). The qualitative results in figures 4 and 5 show that the method successfully reconstructs large furniture pieces and arranges them in a good layout. However, several objects such as plants and chandeliers are disregarded. As the implementation of USL [17] is not publicly available, we only compare our visual results on the Hypersim dataset [52] in figure 7. Similar to InstPIFu, USL misses many objects, and the reconstructed components have a simplified geometry with no texture. In contrast, our results are much more realistic and have higher visual quality.
Additionally, while we use DreamGaussian [58] as part of our framework, we also compare with using DreamGaussian by itself in a non-compositional approach. To apply the model, we treat all the instances in a scene as a single object and reconstruct them together. As the model has seen several scenes composed of more than one object during its training [11], it performs reasonably under this setting. Given the different camera intrinsics in the evaluation, we also compare the results of applying the model on the images after using a reprojection similar to the one used in our pipeline. This further boosts its performance as measured in table 2. Still, its results are of significantly lower quality compared to our compositional approach, which can be analyzed in figure 6.
Lastly, we qualitatively evaluate in figure 8 the performance of InstPIFu [37], DreamGaussian [58] and our method on real-world data with diverse scenarios. The results show that the proposed solution reconstructs complex scenes well, while overcoming many limitations of prior works.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Input image DreamGaussian[58] Ours Ground Truth
Figure 6: Qualitative results on HOPE-Image [61].
Refer to caption Refer to caption Refer to caption Refer to caption
Input image USL [17] Ours (input view) Ours (another view)
Figure 7: Qualitative results on the Hypersim dataset [52]. We compare against the USL method which also does not require 3D data for supervision. Our scenes are more complete, with significantly more objects being recognized and reconstructed.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Input image InstPIFu [37] DreamGaussian [58] Ours
Figure 8: Qualitative results on real-world data.
4.3 Ablations
The two most important interfaces in combining the depth estimation and the single-view object reconstruction components are our reprojection and amodal completion. We present an ablation of these components in table 4. The object reconstruction module $\mathcal{R}$ expects an image of an object centered in the camera frustum. Therefore, simply feeding an image crop from different parts of the scene image will lead to unrealistic views and, consequently, deformed output objects. Additionally, we use amodal completion to contend with the various occlusions that appear in the input view. This step significantly boosts the performance on the 3D-FRONT dataset, but does not improve the numerical evaluation on HOPE-Image dataset, where most of the objects in the scenes are not occluded.
3D-FRONT [15] HOPE-Image [61]
Method Chamfer $\downarrow$ F-Score $\uparrow$ Chamfer $\downarrow$ F-Score $\uparrow$
Ours w/o reprojection 0.155 58.56 1.505 51.89
Ours w/o amodal completion 0.122 66.81 1.450 54.99
Ours (full) 0.120 68.82 1.446 54.82
Table 4: Ablation results for our reprojection and amodal completion procedures. Both components contribute to the performance of our method.
4.4 Limitations
The proposed method has certain shortcomings and there is significant room for improvement, especially for in-the-wild predictions. As we rely on the unprojected depth for the layout of the scene, when the camera calibration or the scale and offset estimation fail to match the correct units, the reconstructed scene will have an erroneous structure, with objects colliding with one another.
We also inherit the limitations of the single-view object reconstruction method. In our experiments, we notice that DreamGaussian sometimes overestimates the size of the object along the $z$-dimension, when it cannot be observed in the input view. Such cases are prominent in figure 6, e.g., orange juice box in the last example. Additionally, DreamGaussian requires the estimation of the perceived camera elevation of the input crop with respect to the normalized object pose. As this cannot be robustly inferred, it significantly affects the reconstruction performance of the model. We believe that most of the current limitations can be overcome by improving the implementation of some of the particular modules in our framework and by enhancing their interoperability. For example, conditioning the depth estimation on the inferred camera calibration can help disambiguate depth scale [55]. Also, the 3D object reconstruction module can benefit from having the estimated instance depth as an additional input [22, 62]. Lastly, relying more on the overall scene context during the amodal completion stage can improve the recovery of missing object parts due to occlusion, as it has been recently shown in concurrent works [46, 64].
5 Conclusion
In this work, we introduce a modular framework for reconstructing complex 3D scenes from an image. We prioritize generalization by taking a divide-and-conquer approach as opposed to an end-to-end model. Our decomposing into multiple entities benefits from existing components, which effectively solve the established subtasks. We develop the necessary interfaces that enable the sub-components to function properly and finally yield a full 3D reconstruction. Our experiments decidedly show the advantage of the proposed method on various types of scenes. Considering the illustrated performance on the diverse scenarios, we believe that the proposed method is a strong baseline and a stepping stone towards achieving generalizable full 3D scene reconstruction from a single image.
Acknowledgement
This work was funded by the German Federal Ministry of Education and Research (BMBF), FKZ: 01IS22082 (IRRW). The authors are responsible for the content of this publication. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b112dc IRRW. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683.